OpenStudy (anonymous):

what can i multiply to simplify 30/48

OpenStudy (solomonzelman):

30/48 is reducible by dividing the top and bottom by 6.

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{30}{48}=\frac{30\color{blue}{\div 6}}{48\color{blue}{\div 6}}= \frac{?}{?} }\) can you finish this problem for me please ?